# Noise analysis of a quasi-phase-matched quantum frequency converter and higher-order counter-propagating SPDC

### v1

@@ 3.3 Parametric noise and high-order counter-propagating SPDC
In this section we discuss the different contributions arising from SPDC in our noise measurements. Figure 6 presents a high resolution measurement of the noise spectrum from 1320 nm to 1620 nm. The red component of the curve indicates the region covered with a high-resolution measurement realised via the combination of a narrow-band tunable filter and a superconducting nanowire single-photon detector (SNSPD). The filter has a minimum bandwidth of 38 pm, a minimum stepsize of 1 pm, and a scan range from 1480 nm to 1620 nm. To maintain a feasible measurement time, the measurements were taken with both a bandwidth and step size of 100 pm. The blue portion of the graph was recorded with a InGaAs single-photon spectrometer from 1320 nm to 1490 nm (resolution 0.55 nm).
@@
Figure 6 plots the measured noise background for the configuration where both the polarisation of pump and the detection are aligned with the -axis of the crystal. The noise shows two distinct spectral feature: a slowly varying broad background and a succession of sharp peaks of different heights, which are not perfectly equidistant across the telecommunication wavelengths. We attribute the broad background noise to parasitic SPDC driven by poling errors and confirm this via temperature tuning of the crystal. The temperature tuning of spectral features of the broad noise floor roughly agrees with the prediction from temperature-dependent Sellmeier equations and the thermal expansion of the crystal[40, 41]. See Figure 7 and the Table S2 in the supplemental document for more details on the temperature tuning measurements. The measured generated NSD of the parasitic SPDC background is (1104) kHz/nm for a conversion efficiency of (72.30.4)% at (74.50.3) W circulating pump power at the target wavelength of 1587 nm [29], which corresponds to a noise floor is about 5 times smaller then in the best state-of-the-art single-step converter which is based on a ppLN waveguide [12]. Previous measurements of the phase-matching curve at very large detunings predicted a lower maximum for the noise floor of only 20 kHz/nm at unity conversion efficiency [31]. We attribute this 5-fold discrepancy between the experimentally predicted and measured NSD to the difference in our theoretically anticipated and experimentally realised values of . In the simple model for frequency conversion [20], the NSD attributed to parasitic SPDC for unity conversion efficiency should be independent of both the pump-power and the effective non-linearity. However, the converter efficiency deviated from its theoretically predicted dependence on pump power, with an inferred of 177 W compared to a 2 to 3 times lower theoretically predicted . The measured NSD shows the expected linear dependence on pump power. One potential reason for the relatively large deviation between the predicted and inferred is a combination of effects not contained in the simple modeling, which does not fully incorporate imperfect mode-matching of the input beam, angle-anisotropy of the indices-of-refraction for the pump, input and output wavelengths and the spatially multi-mode nature of the input and output fields in the regime of high conversion efficiencies (input depletion), effect of which we in fact observed in the mode-shape of the depleted input light after conversion at high efficiency.
@@
Around the absorption edge of the crystal (idler wavelengths of 4.15 m with a corresponding signal wavelength of 1430 nm) we observe a increase in the SPDC noise floor which then plateaus towards shorter signal (longer idler) wavelengths. One potential explanation for this increase in the noise floor is the partial absorption of the idler mode reducing the effective crystal length and therefore degrading the coherent suppression of highly phase-mismatched parasitic SPDC. In the literature the absorption edge of KTP is suggested to be between 4000 nm to 4500 nm [42] and likely varies between different methods of crystal growth.
@@
The series of narrow-band peaks are identified to be higher-order counter-propagating SPDC processes. These peaks emerge when the design poling period matches an integer multiple of the theoretical poling period for the counter-propagating process. As a result, these peaks lie equidistant in the wavevector mismatch . A detailed theoretical background on higher-order counter-propagating SPDC is given in Section 4. Figure 6 plots the relevant measured spectra, with the order m of the processes denoted on the axes above the plot, labelling all four SPDC processes for all possible directions the signal and idler photons can be emitted into relative to the pump beam. The phase-matching order is predicted from the ratio , where is the grating vector and the poling period. The "" indicates the direction in which the signal photons and the idler photons are travelling relative to the pump beam. Note that the quasi-phase-matching order for counter-propagating signal (green) increases with increasing signal wavelength and vice-versa for the idler (pink). This is because the corresponding order necessary to satisfy phase-matching with the fixed grating will increase (decrease) with the increasing (decreasing) the energy of the counter-propagating field.
In an ideal periodically-poled crystal with 50:50 duty cycle, only odd orders of m are phase-matched. Even-order higher-order peaks are present due to poling errors in the duty cycle.
@@
Because we utilise a standing-wave cavity for the pump field, we collect signal photons corresponding to both counter-propagating processes in our measurements. For the process where the idler photon counter-propagates orders from to are visible. For the alternate process, where the signal photon counter-propagates relative to the pump field, we observe orders from to . As we move towards shorter signal wavelengths and accordingly longer idler wavelengths that lie beyond the transparent region of the crystal, we observed a broadening of the narrow peaks that we attribute to partial absorption of the idler field. This is especially evident for the 11th and 43th order peaks. Further, in this region where the crystal becomes increasingly opaque for the idler light, the accuracy of the Sellmeier equations [40] become limited and accordingly so to our prediction for the phase-matching wavelengths.
@@
To our knowledge, the observation of such higher-order counter-propagating peaks was first evident in the measurements of in Ref. [43] in ppLN waveguides, though absent an explanation of their origin. In Ref. [24] the narrow-band peaks are not evident but may have been overwhelmed by the comparatively strong broadband noise. This is a noise source that should arise in all frequency converters based on quasi-phase matched sources where the pump lies between the interconnected wavelengths. Notably, in the limit of an ideal quasi-phase-matched crystal - which would ideally suppress the contribution of parasitic phase-matching that plagues current implementations - these noise peaks are most pronounced. Owing however, to the narrow bandwidths of these features, with luck they will likely lie outside the design bandwidth of a converter.
@@
All narrow peaks (even and odd orders) are substantially broadened up by about a factor of 7-35 – not only in the region where the idler experiences increased absorption. Compare Table S3 and S4 in the supplemental document. In our preliminary measurements we investigated two crystals with identical specifications from different manufacturers (Raicol and Tailored Photons). The measurements presented here were made with the Raicol crystal, as it was the basis of the converter characterisation of Ref. [29] and thus provides an accurate normalisation for the NSD. Notably, while the two crystals showed qualitatively similar noise spectra, there were a few distinct differences. The peaks of the counter-propagating processes evident in the spectrum taken with the Raicol crystal are considerably broader then those of the Tailored Photons (TP) crystal. Further, the process where the signal counter-propagates is only evident in measurements of the Raicol crystal. Also note that the background noise level due to the ordinary noise process is about 5 times higher for the Tailored Photons crystal compared to the Raicol crystal. These distinctions might suggest different magnitudes for different types of poling errors.
@@
Our interpretation of the measured background is further supported by measurements of the temperature dependence of the NSD. As shown in Figure 7, the broad SPDC background demonstrates a clear temperature dependence expected from a co-propagating process. In contrast, the narrow peaks demonstrate the comparatively weak temperature dependence expected from counter-propagating processes.
@@
Knowledge of these counter-propagating noise processes might provide valuable insight into the underlying quasi-phase matching, or the crystal properties. For example, with good knowledge of the poling period, and the Sellmeier equations for the signal wavelength, the location of the peaks at the signal wavelength allow inference of the Sellmeier equations at the idler wavelength. Alternatively, the periodic nature of these features might also allow for improved precision in inferring the grating period. Furthermore, the appearance or absence of phase-matching associated higher-order polings may complement our understanding of poling errors, especially when considered alongside the size of the homogeneous-background attributed to random duty-cycle errors.
@@ 4 Theory of higher-order counter-propagating SPDC
Here, is the pump field, L the length of the crystal, the nonlinear coefficient, and D the duty cycle of the periodic poling. Note that the efficiency of higher-order processes scales with order as . Thus, for large the efficiency decreases relatively slowly with increasing order . Due to the term in the above equation, even for small deviations of D from 50:50 (e.g. D=0.49), both odd and even orders of large m (e.g. m<10) can be observed in the spectrum. The phase-mismatch of a SPDC process, taking into account -th order quasi-phase-matching, is given by
@@
Here was used. Thus, counter-propagating processes detune very fast in signal frequency compared to the ordinary processes. This can also be seen when comparing the colored axes above Figure 6: For the ordinary processes not even one order is crossed over, over the whole measurement interval, while for the counter-propagating processes 9 orders are crossed over. This also leads to different bandwidths in frequency - these highly non-degenerate processes are approximated in first order by
Here are the group indices. The "-"-sign is for ordinary and the "+"-sign for counter-propagating processes. Note that the counter-propagating processes are about 100 times more narrow than the ordinary processes and are thus apparent as very narrow peaks in the spectrum. The bandwidths of the relevant processes are that of the converter 109 GHz (0.92 nm @1587 nm), the ordinary processes with 354 GHz (2.97 nm @1587 nm) and the counter-propagating processes with 3.6 GHz (30 pm @1587 nm). When detuning the crystal temperature by we have
Thus, ordinary processes detune fast with temperature and counter-propagating processes detune slow. For counter-propagating processes a substantial contribution of the temperature tuning (about half) is due to the temperature dependence of the grating vector . The temperature dependence of the grating vector is due to the thermal expansion of the poled domains. Since the grating vector is multiplied by m, for bigger m, this contribution increases.
@@
Thus the fraction equals the phase-matching order m at the peaks (compare the four colored axes in Figure 6). These fractions are plotted as additional axes above Figure 6. From the distance of the peaks relative to each other, the grating period can be calculated
This could be used to calibrate Sellmeier equations and effective poling periods, e.g already a 0.1% different poling length would shift the sharp peaks measurably.
5 Discussion
In this work we have provided a full characterisation of the noise spectrum from 1140 nm to 1650 nm of a bulk ppKTP quantum frequency converter pumped at 1064 nm, making use of measurements over a very broad spectral range combining both a large dynamic range and high spectral resolution. Through selection of the polarisation of the pump and detected light, as well as tuning of the crystal temperature, different peaks associated with Stokes-Raman scattering were identified in the range from 1140 nm to 1330 nm (up to 60 THz from the pump) by comparing the noise spectrum to previous Raman scattering measurements in KTP [35, 36, 37]. At photon energies 60 THz or lower than that of the pump light – between 1330 nm to 1650 nm – the noise was found to be dominated by broadband parasitic SPDC and phase-matched high-order counter-propagating SPDC.
Higher-order counter-propagating SPDC processes were observed in the spectrum in the form of a characteristic succession of very narrow-band peaks of different height. We believe this phenomenon was heretofore unobserved experimentally. Additionally, phase-matching of even orders of m, enabled by small poling errors related to the duty-cycle were also visible.
As these counter-propagating SPDC processes are very sensitive to the material dispersion of the host nonlinear crystal, such measurements could be used to refine Sellmeier equations or to precisely measure the poling period. Thus our findings could potentially be useful for example to refine the mid-IR Sellmeier equations. Also the presented methodology may give additional, so far inaccessible information on the specifics of poling errors and their distribution, e.g. by the existence, the relative brightness and the shape of even and odd order peaks.
@@ 5 Discussion
Acknowledgments We thank Wiliam Staunton and Oliver Benson for providing equipment.

### v2

@@ 3.3 Parametric noise and high-order counter-propagating SPDC
In this section we discuss the different contributions arising from SPDC in our noise measurements. Fig. 6 presents a high resolution measurement of the noise spectrum from 1320 nm to 1620 nm. The red component of the curve indicates the region covered with a high-resolution measurement realised via the combination of a narrow-band tunable filter and a superconducting nanowire single-photon detector (SNSPD). The filter has a minimum bandwidth of 38 pm, a minimum step size of 1 pm, and a scan range from 1480 nm to 1620 nm. To maintain a feasible measurement time, the measurements were taken with both a bandwidth and step size of 100 pm. The blue portion of the graph was recorded with a InGaAs single-photon spectrometer from 1320 nm to 1490 nm (resolution 0.55 nm).
@@
Fig. 6 plots the measured noise background for the configuration where both the polarisation of pump and the detection are aligned with the -axis of the crystal. The noise shows two distinct spectral feature: a slowly varying broad background and a succession of sharp peaks of different heights, which are not perfectly equidistant across the telecommunication wavelengths. We attribute the broad background noise to parasitic SPDC driven by poling errors and confirm this via temperature tuning of the crystal. The temperature tuning of spectral features of the broad noise floor roughly agrees with the prediction from temperature-dependent Sellmeier equations and the thermal expansion of the crystal[40, 41]. See Fig. 7 and Table S2 in the supplemental document for more details on the temperature tuning measurements. For a circulating pump power of W, we achieve our highest measured conversion efficiency of % while generating kHz/nm noise counts at the target wavelength of 1587 nm [29]. This corresponds to a noise floor that is about 5 times smaller than that of the best state-of-the-art single-step converter which is based on a ppLN waveguide [12] and about 20 times larger than the unpoled bulk converter presented in Ref. [28]. Previous measurements of the phase-matching curve at very large detunings predicted a lower maximum for the noise floor of only 20 kHz/nm at unity conversion efficiency [31]. We attribute this 5-fold discrepancy between the experimentally predicted and measured NSD to the difference in our theoretically anticipated and experimentally realised values of . In the simple model for frequency conversion [20], the NSD attributed to parasitic SPDC for unity conversion efficiency should be independent of both the pump power and the effective non-linearity. However, the converter efficiency deviated from its theoretically predicted dependence on pump power, with the experimentally inferred of 177 W being 2 to 3 times higher than that theoretically predicted. The measured NSD shows the expected linear dependence on pump power. One potential reason for the relatively large deviation between the predicted and inferred is a combination of effects not contained in the simple modelling, which does not fully incorporate imperfect mode-matching of the input beam, angle-anisotropy of the indices-of-refraction for the pump, input and output wavelengths and the spatially multi-mode nature of the input and output fields in the regime of high conversion efficiencies (input depletion) - the latter being an effect we observe in the mode-shape of the depleted input light after conversion at high efficiency.
@@
Around the absorption edge of the crystal (idler wavelengths of 4.15 m with a corresponding signal wavelength of 1430 nm) we observe an increase in the SPDC noise floor which then plateaus towards shorter signal (longer idler) wavelengths. One potential explanation for this increase in the noise floor is the partial absorption of the idler mode reducing the effective crystal length and therefore degrading the coherent suppression of highly phase-mismatched parasitic SPDC. In the literature the absorption edge of KTP is suggested to be between 4000 nm to 4500 nm [42] and likely varies between different methods of crystal growth.
@@
The series of narrow-band peaks are attributed to higher-order counter-propagating SPDC processes. These peaks emerge when the design poling period matches an integer multiple of the theoretical poling period for the counter-propagating process. As a result, these peaks lie equidistant in the wavevector mismatch . A detailed theoretical background on higher-order counter-propagating SPDC is given in Section 4. Fig. 6 plots the relevant measured spectra, with the order m of the processes denoted on the two axes above the plot for the two counter-propagating SPDC processes. The phase-matching order is predicted from the ratio , where is the grating vector and the poling period. The "" indicates the direction in which the signal photons and the idler photons are travelling relative to the pump beam. Note that the quasi-phase-matching order for the counter-propagating signal (green) increases with increasing signal wavelength and vice-versa for the idler (pink). This is because the corresponding order necessary to satisfy phase-matching with the fixed grating will increase (decrease) with the increasing (decreasing) energy of the counter-propagating field. While for an idealised periodically-poled crystal with 50:50 duty cycle only odd orders of m are phase-matched, here the presence of poling errors in the duty cycle allows for phase-matching of the even orders. This is a consequence of the higher-order phase-matching, where an effective net contribution to the SPDC amplitude is only made in the last portion of each domain. Consequently, a deviation of only a few percent in the design duty cycle can amount to an appreciable deviation for a higher-order process where only the last few microns of each domain count.
@@
Because we utilise a standing-wave cavity for the pump field, we collect signal photons corresponding to both counter-propagating processes in our measurements. For the process where the idler photon counter-propagates orders from to are visible in our measurements. For the alternate process, where the signal photon counter-propagates relative to the pump field, we observe orders from to . As we move towards shorter signal wavelengths and accordingly longer idler wavelengths that lie beyond the transparent region of the crystal, we observe a broadening of the narrow peaks that we attribute to partial absorption of the idler field. This is especially evident for the 11th and 43th order peaks. Further, in this region where the crystal becomes increasingly opaque for the idler light, the accuracy of the Sellmeier equations [40] becomes limited and accordingly so to do our predictions for the exact phase-matching wavelengths.
@@
To our knowledge, the observation of such higher-order counter-propagating peaks was first evident in the measurements of in Ref. [43] in ppLN waveguides, though absent an explanation of their origin. In Ref. [24] the narrow-band peaks are not evident but may have been overwhelmed by the comparatively strong broadband noise. This is a noise source that should arise in all frequency converters based on quasi-phase-matched sources where the pump lies between the interconnected wavelengths. Notably, in the limit of an ideal quasi-phase-matched crystal - which would ideally suppress the contribution of parasitic phase-matching that plagues current implementations - these noise peaks are most pronounced - albeit only for odd orders. Owing however, to the narrow bandwidths of these features, they will most often lie outside the design bandwidth of a converter.
@@
All narrow peaks (even and odd orders) are substantially broadened up by about a factor of 7-35 – not only in the region where the idler experiences increased absorption. Compare Table S3 and S4 in the supplemental document. In our preliminary measurements we investigated two crystals with identical specifications from different manufacturers (Raicol and Tailored Photons). The measurements presented here were made with the Raicol crystal, as it was the basis of the converter characterisation of Ref. [29] and thus provides an accurate normalisation for the NSD. Notably, while the two crystals showed qualitatively similar noise spectra, there were a few distinct differences. The process where the signal counter-propagates is only evident in measurements of the Raicol crystal. We attribute this to the Tailored Photons (TP) crystal having an approximately 5 times larger background for the ordinary noise process when compared to the Raicol crystal, effectively burying the higher-order process. This suggests the random duty cycle errors in the TP crystal are larger than those present in the Raicol crystal. However, despite this reduced background, the peaks of the counter-propagating processes evident in the spectrum taken with the Raicol crystal are considerably broader than those of the TP crystal. This suggests the Raicol crystal might have comparatively larger poling length errors, that, while evident in this measurement, are not deleterious to the SPDC noise floor.
@@
Our interpretation of the measured background is further supported by measurements of the temperature dependence of the NSD. As shown in Fig. 7, the broad SPDC background demonstrates a clear temperature dependence expected from a co-propagating process. In contrast, the narrow peaks demonstrate the comparatively weak temperature dependence expected from counter-propagating processes.
@@
Knowledge of these counter-propagating noise processes might provide valuable insight into the underlying quasi-phase-matching, or the crystal properties. For example, with good knowledge of the poling period, and the Sellmeier equations for the signal wavelength, the location of the peaks at the signal wavelength allows inference of the Sellmeier equations at the idler wavelength. Alternatively, the periodic nature of these features might also allow for improved precision in inferring the grating period. Furthermore, the appearance or absence of phase-matching associated higher-order polings may complement our understanding of poling errors, especially when considered alongside the size of the homogeneous-background attributed to random duty cycle errors.
@@ 4 Theory of higher-order counter-propagating SPDC
Here, is the pump field, L the length of the crystal, the nonlinear coefficient, and D the duty cycle of the periodic poling. Note that the efficiency of higher-order processes scales with order as . Thus, for large the efficiency decreases relatively slowly with increasing order . Due to the term in the above equation, even for small deviations of D from 50:50 (e.g. D=0.49), both odd and even orders of large m (e.g. m>10) can be observed in the spectrum. The phase-mismatch of a SPDC process, taking into account -th order quasi-phase-matching, is given by
@@
Here was used. Thus, counter-propagating processes detune very fast in signal frequency compared to the ordinary processes. This also leads to different bandwidths in frequency - these highly non-degenerate processes are approximated in first order by
Here are the group indices. The "-"-sign is for ordinary and the "+"-sign for counter-propagating processes. Note that the counter-propagating processes are about 100 times more narrow than the ordinary processes and are thus apparent as very narrow peaks in the spectrum. The bandwidths of the relevant processes are that of the converter: 109 GHz (0.92 nm at 1587 nm), the ordinary processes: 354 GHz (2.97 nm at 1587 nm) and the counter-propagating processes: 3.6 GHz (30 pm at 1587 nm). When detuning the crystal temperature by we have
Thus, ordinary processes detune fast with temperature and counter-propagating processes detune slowly. For counter-propagating processes, a substantial contribution of the temperature tuning (about half) is due to the temperature dependence of the grating vector . The temperature dependence of the grating vector is due to the thermal expansion of the poled domains. Since the grating vector is multiplied by m, for bigger m, this contribution increases.
@@
Thus the fraction equals the phase-matching order m at the peaks (compare the four colored axes in Fig. 6). These fractions are plotted as additional axes above Fig. 6. From the distance of the peaks relative to each other, the grating period can be calculated
This could be used to calibrate Sellmeier equations and effective poling periods, e.g. already a 0.1% different poling length would shift the sharp peaks measurably.
5 Conclusion
In this work we have provided a full characterisation of the noise spectrum from 1140 nm to 1650 nm of a bulk ppKTP quantum frequency converter pumped at 1064 nm, making use of measurements over a very broad spectral range combining both a large dynamic range and high spectral resolution. Through selection of the polarisation of the pump and detected light, as well as tuning of the crystal temperature, different peaks associated with Stokes-Raman scattering were identified in the range from 1140 nm to 1330 nm (up to 60 THz from the pump) by comparing the noise spectrum to previous Raman scattering measurements in KTP [35, 36, 37]. At photon energies 60 THz or lower than that of the pump light – between 1330 nm to 1650 nm – the noise was found to be dominated by broadband parasitic SPDC and phase-matched high-order counter-propagating SPDC processes.
These higher-order counter-propagating SPDC processes were observed in the spectrum in the form of a characteristic succession of very narrow-band peaks of different heights. We believe this phenomenon was heretofore unobserved experimentally. Additionally, phase-matching of even orders of m, enabled by small poling errors related to the duty cycle was also visible.
As these counter-propagating SPDC processes are very sensitive to the material dispersion of the host nonlinear crystal, they may find application in refining Sellmeier equations or the precise measurement of gratings. The enhanced sensitivity of these counter-propagating SPDC processes to poling errors also suggests they could be exploited to provide insights into these errors and their distribution, e.g. through the existence or absence, the relative brightness and the widths of peaks. Our presented measurement techniques - in essence combining high power pumping at 1064 nm and measurements with off-the shelf InGaAs-grating spectrometers - both of which are technologically rather straight-forward and moderately cost-effective - might offer another approach to better understand and design improved periodically poled nonlinear devices.
@@ 5 Conclusion
Acknowledgments We thank William Staunton and Oliver Benson for providing equipment.